# Partially-Supervised Neural Network Model For Quadratic Multiparametric Programming

### v1

@@ Abstract
keywords:
Partially-Supervised Neural Network , Analytical Derivation of Model Parameters , Multiparametric Programming , Neural Network for Optimization
journal: Computers and Operations Reseerch
[…]
organization=Management Science and Engineering, University of Waterloo,addressline=200 University Ave W, city=Waterloo, postcode=N2L 3G1, state=ON, country=Canada
{highlights}
We propose a NN model that exactly represents the solution function of QP.
We explain inaccuracies of traditional black-box NN-based QP solvers.
We derive some model parameters directly from the problem coefficients.
Proof of concept support for the method is provided using empirical datasets.
The model produces solutions instantly with less than 1E-10 squared KKT violations.
@@ 1 Introduction
There has been increasing interest in using Neural Networks (NN) to predict the solutions to complex nonlinear optimization problems in energy management, chemistry, control theory, and other domains [1, 2, 3]. Effectively, this goal amounts to estimating a global solution function [^1] The function mapping parameters to optimal primal solutions is sometimes referred to as an optimizer in the multiparametric programming literature. We use the term solution function to refer to the mapping between parameters to both primal and dual solutions. that predicts optimal solutions based on a given set of input parameters, such as the right hand side (RHS) vector and cost coefficients of an optimization model. Most applications treat the NN model as a black-box and employ standard training methods using datasets of input-output pairs, representing particular values of system parameters and corresponding optimal solutions. Most also ignore the mathematical structure of the underlying function and utilize a generic deep NN (DNN) architecture to approximate solutions. Such modeling approaches necessitate large, computationally expensive training datasets to achieve satisfactory performance, yet cannot guarantee the feasibility or optimality of results.
The multiparametric programming literature similarly seeks to represent the global solution function to optimization problems by characterizing the optimal solutions as a function of feasible problem parameters, and identifying the regions of parameter space where these functions are valid [4]. Recent work has attempted to integrate multiparametric programming with DNNs, based on recognition of their similar underlying mathematical characteristics [2, 3]. For multiparametric Linear Programming (mp-LP) and Quadratic Programming (mp-QP) with linear constraints, the optimal primal and dual solutions are piecewise linear (PWL) functions of the parameters [4]. Specifically, each set of constraints binding at a solution corresponds with a subregion of the feasible domain, called critical region, where solutions change linearly with respect to changes in the RHS parameters, and solutions change piecewise linearly between adjacent critical regions. Similarly, a NN with ReLU activation functions is a general form of PWL function with trainable weights and biases [5]. In theory, therefore, for any mp-LP or mp-QP, one should be able to find an optimal set of weights and biases such that a NN can exactly represent the corresponding solution function [2].
@@ 2.1 Surrogate Models
Using NNs for surrogate modeling of optimization problems is an established field with various applications in engineering such as control [2] and chemistry [3], but it is largely studied in the literature on power systems. The approaches leverage DNNs as universal function approximators to estimate the nonlinear relationship between problem parameters and optimal solutions, which has found success in power management [1, 6, 7]. Despite their advantages as nonlinear solvers, guaranteeing feasibility and optimality is challenging, and finding the best model parameters from a large parameter space is data-greedy with significant computational training burdens.
To improve feasibility, [8] and [9] proposed two-stage models that first predict optimal solutions with a NN and then post-process the solutions to enforce feasibility. [7] proposed a NN architecture that enforces optimal dispatch generator limits using a sigmoid activation function in the output layer. [6] processed optimal solutions through sequentially connected sub-networks reflecting problem characteristics. Physics-informed Neural Networks (PINNs) use DNNs that are also trained to satisfy KKT optimality conditions, resulting in reduced dataset requirements, more accurate predictions with less violations, and more generalizability power [1].
Other studies in the OPF literature are concerned with using ML to improve execution time of solvers. In a series of studies (e.g. [10, 11]), researchers observed that it is possible to reduce the dimensions of the problem by predicting the active constraints. Specifically, the solution must always satisfy the equality constraints and predicting the binding inequality constraints reduces the size of the problem space. For this purpose, different types of classifiers were trained to find the solution faster [10, 11, 12]. Others [13] proposed faster solutions to the AC-OPF problem by removing inactive constraints to modify the feasibility space and downsize the problem.
@@ 2.2 Multiparametric Programming
Multiparametric programming became more popular in the 2000s when researchers showed that model predictive control laws can be expressed as multiparametric programming solutions [14], specifically in the form of a piecewise linear function of the parameter set. Studies proposed a geometric interpretation that considers the critical regions of mp-LP solutions as polyhedra, where the parameter space is explored by moving the parameter set among critical regions [15]. However, it was later shown that this approach does not guarantee discovery of all critical regions [16]. The geometric approach was later generalized to solve mp-QP by Bemporad et al. [17] as it shares similar characteristics with mp-LP. In contrast to mp-LP, which relies on the simplex algorithm, multiparametric QP was theorized using the Basic Sensitivity Theorem proposed by Fiacco [18]. It was shown that solutions to mp-QP are PWL functions and the optimal objective function is a piecewise quadratic function of the parameter set [14, 4].
@@
The geometric approach involves finding critical regions by solving the problem for an initial parameter set, $\boldsymbol{\theta}_{0}$, identifying active sets and obtaining the parametric solution corresponding to the critical region where the solution is valid. Subsequently, the remaining critical regions are explored by moving outside the initial critical region. The original study by [14] proposes constraint reversal by successively reversing signs of each constraint, i.e., from $A_{i}x\leq b_{i}$ to $A_{i}x>b_{i}$. However, this approach creates too many artificial cuts leading to scalability challenges for larger problems. Two main modifications were proposed to overcome scalability. First is an active set inference approach [16] that adds or drops one constraint at a time from the active constraint set to move to an adjacent region. The second is a variable step size approach [19] that chooses the facets defining the critical region, finding its centre and moving outside to discover an adjacent critical region. However, both methods fail to explore the entire feasible parameter set when facet-to-facet property does not hold.
@@
The critical regions of an mp-QP problem are each defined by a unique corresponding set of active constraints, so the feasible space can be exhaustively explored by enumerating all combinations of constraints and removing infeasible combinations. However, as the number of constraint combinations grows exponentially, this approach does not scale up to higher dimensional problems. The number of potential combinations can be reduced when an infeasible set is discovered [20] since a power of an infeasible set is also infeasible, so branch and bound type algorithms have been proposed to discover the critical regions [21, 22].
@@
The mixed integer versions of LP and QP problems can be viewed as combinations of multiple mp-LP and mp-QP problems. Therefore, strategies proposed to solve mp-MILP and mp-MIQP problems include enumerating all combinations of integer variables, solving the corresponding LP or QP problems, and combining solutions by comparison. To make the solution tractable, researchers again proposed branch and bound [23] and decomposition strategies [24].
@@
The properties of critical regions are different for mp-MIQP and mp-MILP. Whereas the critical regions for mp-MILP are polytopes, obtaining optimal solutions by comparing multiple mp-QP solutions can result in quadratically constrained critical regions. Maintaining critical regions in the form of polyhedra offers computational advantages, so researchers have applied McCormick relaxations for linearizing the objective function or critical regions [25]. Other applications of multiparametric programming to optimization, including multiparametric nonlinear programming, are outside the scope of this study (see [26] for further review.)
@@
Examining mathematical properties of mp-LP and mp-QP problems under a multiparametric setting led researchers to recognize their similarity to the PWL characteristics of DNNs. [3] attempted to integrate PWL behaviour of the problem solution analytically. [2] examined the functional similarities between DNNs and model predictive control (MPC) laws, and trained a model to estimate MPC solutions for a time variant system. Similarly, [27] proposed an approach to integrate MPC laws of a MILP problem for a microgrid system. However, these approaches employ black-box models that do not exactly represent the solution function. Moreover, none of these works compare optimality and feasibility between their models and traditional solver benchmarks or DNN based approaches. This paper proposes a custom NN architecture that closely aligns with the PWL structure of the global solution function to improve accuracy and generalizability of mp-QP predictions.
@@ 3.1 Neural Networks
Neural Networks are universal function approximators that are defined as composite functions that can be expressed as an indirect mapping between inputs x and outputs y, $f:\textbf{x}\rightarrow\textbf{y}$. The basic type of NN, known as shallow feed forward neural network, consists of an input and output layer, connected by a hidden layer, $h$. Loosely influenced by neurons and synapses of the brain, the value of each neuron is determined by the weighted sum of the previous neurons’ values, which is called input current denoted by $z_{i}$ for node $i$. However, the neuron’s value is not always defined as its input current but usually transformed by an activation function. For example, the rectified linear unit (ReLU) activation function is a threshold function that is activated only if the input current value is positive, expressed mathematically as follows:
@@
For a shallow model where $n_{h}=1$, $\textbf{W}^{1}\in\mathbb{R}^{d_{x}\times d_{h}}$, $\textbf{b}^{1}\in\mathbb{R}^{d_{h}}$,$\textbf{W}^{2}\in\mathbb{R}^{d_{h}\times d_{y}}$ and $\textbf{b}^{2}\in\mathbb{R}^{d_{y}}$.
@@ 4.2 Adding Inequality Constraints
As discussed earlier, when inequality constraints are binding, the active constraints can be considered as equality constraints with their the optimal dual variables, $\boldsymbol{\mu}^{*}>0$. Then, the solution can be found by expanding the Lagrangian function and finding its critical points.
@@ 4.3 Predicting Active Constraints and Shadow Prices
For the mp-QP problem 6, $\Theta_{f}\subseteq\Theta$ is a convex set, the primal solution function x$(\boldsymbol{\theta}):\Theta_{f}\rightarrow\mathbb{R}^{n}$ is continuous and piecewise affine. Also the optimal objective function $\textbf{z}(\boldsymbol{\theta}):\Theta_{f}\rightarrow\mathbb{R}$ is continuous and piecewise quadratic.
Proof: The reader can find the proof of the theorem in [4].
@@
Notice that the last row of the above inverse Jacobian matrix, $\textbf{J}^{-1}_{\mathcal{B}_{i}}$, is the gradient vector $\nabla\boldsymbol{\mu}^{*}_{\mathcal{B}_{i}}$ that contains the coefficients of the linear segment of the target PWL function when a certain set of constraints, $\mathcal{B}_{i}$, are binding. In other words, the vector is the set of coefficients that maps $[\textbf{C},\textbf{b}_{e},\textbf{b}_{\mathcal{B}_{i}}]$ and $\boldsymbol{\theta}$ to $\boldsymbol{\mu}^{*}_{\mathcal{B}_{i}}$.
@@ 4.4 Identification of Critical Regions and Data Generation
As explained in the Introduction, the problem characteristics in many mp-QP applications imply that most constraint combinations can never occur, so some method of filtering out the unused combinations is needed. For example, in the above 6-bus DC-OPF case with 6 inequality constraints, the set of all constraint combinations is $\{\{0\},\{1\},\{2\},\{0,1\},\dots,\{0,1,2,3,4,5\}\}$ with $642^{6}=$ 64 elements, yet 59 of these combinations never occur. In the power demand applications investigated in the present study, we use a commercial solver to identify the active constraint combinations by solving the problems over a range of input parameters.
Algorithm 1 details this discovery step. The algorithm first generates data starting from an initial parameter, $\boldsymbol{\theta}^{0}$ and choosing an arbitrary dimension, $k$, and incrementally increasing one dimension from zero to the highest possible feasible value. Then, after solving for each input demand, the set of potential binding constraints is identified by checking whether the associated $\mu_{i}$ is greater than zero for any of the inputs. Then, the slope of the linear segment of the solution function, $\nabla\boldsymbol{\mu}_{\mathcal{B}_{i}}$ is derived following (13). Using a solver to find potential binding constraints also provides insight into the boundaries of critical regions, $[\textbf{CR}_{i}^{-},\textbf{CR}_{i}^{+}]$ corresponding to $\mathcal{B}_{i}$ that are active at solution. This information is used in Algorithm 2 to populate data without using solvers, by sampling random inputs from $[\textbf{CR}_{i}^{-},\textbf{CR}_{i}^{+}]$ for all $i$ and calculating the corresponding optimal solutions using the slope corresponding to each critical region.
@@
Input: Initial point $\boldsymbol{\theta}^{0}$, increment size $\alpha$ and upper limit $\theta^{+}$
@@
Solve problem for all input parameters:
Initiate $i=1$, $\textbf{W}^{0}=[\ ]$, $\textbf{CR}^{-}=[\ ]$, $\textbf{CR}^{+}=[\ ]$
@@
     $\textbf{CR}_{i}^{-}=\boldsymbol{\theta}$
     (i) Solve problem for $\boldsymbol{\theta}$ using a solver
@@
     (iii) Obtain $(\boldsymbol{\theta},\boldsymbol{\mu}^{*})$ pairs $\nabla\boldsymbol{\mu}^{*}_{\mathcal{B}_{i}}$
     if $\mathcal{B}_{i}\notin\textbf{W}^{0}$ then
         Expand coefficient matrix to calculate $\nabla\boldsymbol{\mu}_{\mathcal{B}_{i}}$ and append to $\textbf{W}^{0}$
         Set $\textbf{CR}_{i}^{+}=\boldsymbol{\theta}$
         Set $i=i+1$
@@
return $\nabla\boldsymbol{\mu}^{*}_{\mathcal{B}_{i}},\textbf{CR}^{-},\textbf{CR}^{+}$

### v2

# Partially-Supervised Neural Network Model For Multiparametric Quadratic Programming

@@ Abstract
Keywords Partially-Supervised Neural Network $\cdot$ Analytical Derivation of Model Parameters $\cdot$ Multiparametric Programming $\cdot$ Neural Network for Optimization
@@ 1 Introduction
There has been increasing interest in using Neural Networks (NN) to predict the solutions to complex nonlinear optimization problems in energy management, chemistry, control theory, and other domains (nellikkath2022physics, ; karg2020efficient, ; katz2020integrating, ). Effectively, this goal amounts to estimating a solution function that predicts optimal solutions based on a given set of input parameters, such as the right-hand-side (RHS) vector and cost coefficients of an optimization model. Most applications treat the NN model as a black-box and employ standard training methods using datasets of input-output pairs, representing particular values of system parameters and corresponding optimal solutions. Most also ignore the mathematical structure of the underlying function and utilize a generic deep NN (DNN) architecture to approximate solutions. Such modeling approaches necessitate large, computationally expensive training datasets to achieve satisfactory performance, yet cannot guarantee the feasibility or optimality of results.
The multiparametric programming literature similarly seeks to represent the solution function to optimization problems, using algebraic methods to characterize optimal solutions as a function of feasible problem parameters [^1] The function mapping parameters to optimal primal solutions is sometimes referred to as an optimizer in the multiparametric programming literature. We use the term solution function to refer to the mapping between parameters to both primal and dual solutions. and to identify the regions of parameter space where these functions are valid (pistikopoulos2020multi, ). Recent work has attempted to integrate multiparametric programming with DNNs, based on recognition of their similar underlying mathematical characteristics (karg2020efficient, ; katz2020integrating, ). For multiparametric Linear Programs (mp-LP) and Quadratic Programs (mp-QP) with linear constraints, the optimal primal and dual solutions are piecewise linear (PWL) functions of the parameters (pistikopoulos2020multi, ). Specifically, each set of constraints binding at a solution corresponds with a subregion of the feasible domain, called a critical region, where solutions change linearly with respect to changes in the RHS parameters, and solutions change piecewise linearly between adjacent critical regions. Similarly, a NN with ReLU activation functions is a general form of PWL function with trainable weights and biases (montufar2014number, ). In theory, therefore, for any mp-LP or mp-QP one should be able to find an optimal set of weights and biases such that a NN can exactly represent the corresponding solution function (karg2020efficient, ).
@@ 2.1 Surrogate Models
Using NNs for surrogate modeling of optimization problems is an established field with various applications in engineering such as control karg2020efficient , chemistry katz2020integrating , and power systems management (nellikkath2022physics, ; fioretto2020predicting, ; lotfi2022constraint, ). The approach leverages DNNs as universal function approximators to estimate the nonlinear relationship between problem parameters and optimal solutions. Despite the potential advantages of DNNs as nonlinear solvers, guaranteeing feasibility and optimality is challenging, and finding the best model parameters from a large parameter space is data-greedy with significant computational training burdens.
To improve feasibility, zamzam2020learning and chen2023end proposed two-stage models that first predict optimal solutions with a NN and then post-process the solutions to enforce feasibility. lotfi2022constraint proposed a NN architecture that enforces optimal dispatch generator limits using a sigmoid activation function in the output layer. fioretto2020predicting processed optimal solutions through sequentially connected sub-networks reflecting problem characteristics. Physics-informed Neural Networks (PINNs) use DNNs that are also trained to satisfy KKT optimality conditions, resulting in reduced dataset requirements, more accurate predictions with less violations, and more generalizability power (nellikkath2022physics, ).
Other studies in the OPF literature are concerned with using ML to improve execution time of solvers. In a series of studies (e.g. misra2021learning ; deka2019learning ), researchers observed that it is possible to reduce the dimensions of the problem by predicting the active constraints. Specifically, the solution must always satisfy the equality constraints and predicting the binding inequality constraints reduces the size of the problem space. For this purpose, different types of classifiers were trained to find the solution faster misra2021learning ; deka2019learning ; ng2018statistical . Others baker2018joint proposed faster solutions to the AC-OPF problem by removing inactive constraints to modify the feasibility space and downsize the problem.
@@ 2.2 Multiparametric Programming
Multiparametric programming became more popular in the 2000s when researchers showed that model predictive control laws can be expressed as multiparametric programming solutions bemporad2000explicit , specifically in the form of a piecewise linear function of the parameter set. Studies proposed a geometric interpretation that considers the critical regions of mp-LP solutions as polyhedra, where the parameter space is explored by moving the parameter set among critical regions borrelli2003geometric . However, it was later shown that this approach does not guarantee discovery of all critical regions tondel2003algorithm . The geometric approach was later generalized to solve mp-QP by Bemporad et al. bemporad2002explicit as it shares similar characteristics with mp-LP. In contrast to mp-LP, which relies on the simplex algorithm, multiparametric QP was theorized using the Basic Sensitivity Theorem proposed by Fiacco fiacco1976sensitivity . It was shown that solutions to mp-QP are PWL functions and the optimal objective function is a piecewise quadratic function of the parameter set bemporad2000explicit ; pistikopoulos2020multi .
@@
The geometric approach involves finding critical regions by solving the problem for an initial parameter set, $\boldsymbol{\theta}_{0}$, identifying active sets and obtaining the parametric solution corresponding to the critical region where the solution is valid. Subsequently, the remaining critical regions are explored by moving outside the initial critical region. The original study by bemporad2000explicit proposes constraint reversal by successively reversing signs of each constraint, i.e., from $A_{i}x\leq b_{i}$ to $A_{i}x>b_{i}$. However, this approach creates too many artificial cuts leading to scalability challenges for larger problems. Two main modifications were proposed to overcome scalability. First is an active set inference approach tondel2003algorithm that adds or drops one constraint at a time from the active constraint set to move to an adjacent region. The second is a variable step size approach baotic2003new that chooses the facets defining the critical region, finding its centre and moving outside to discover an adjacent critical region. However, both methods fail to explore the entire feasible parameter set when facet-to-facet property does not hold.
@@
The critical regions of an mp-QP problem are each defined by a unique corresponding set of active constraints, so the feasible space can be exhaustively explored by enumerating all combinations of constraints and removing infeasible combinations. However, as the number of constraint combinations grows exponentially, this approach does not scale up to higher dimensional problems. The number of potential combinations can be reduced when an infeasible set is discovered gupta2011novel since a power of an infeasible set is also infeasible, so branch and bound type algorithms have been proposed to discover the critical regions feller2012combinatorial ; feller2013explicit .
@@
The mixed integer versions of LP and QP problems can be viewed as combinations of multiple mp-LP and mp-QP problems. Therefore, strategies proposed to solve mp-MILP and mp-MIQP problems include enumerating all combinations of integer variables, solving the corresponding LP or QP problems, and combining solutions by comparison. To make the solution tractable, researchers again proposed branch and bound acevedo1999algorithm and decomposition strategies jia2006uncertainty .
@@
The properties of critical regions are different for mp-MIQP and mp-MILP. Whereas the critical regions for mp-MILP are polytopes, obtaining optimal solutions by comparing multiple mp-QP solutions can result in quadratically constrained critical regions. Maintaining critical regions in the form of polyhedra offers computational advantages, so researchers have applied McCormick relaxations for linearizing the objective function or critical regions oberdieck2015explicit . Other applications of multiparametric programming to optimization, including multiparametric nonlinear programming, are outside the scope of this study (see oberdieck2016multi for further review.)
@@
Examining mathematical properties of mp-LP and mp-QP problems under a multiparametric setting led researchers to recognize their similarity to the PWL characteristics of DNNs. katz2020integrating attempted to integrate PWL behaviour of the problem solution analytically. karg2020efficient examined the functional similarities between DNNs and model predictive control (MPC) laws, and trained a model to estimate MPC solutions for a time variant system. Similarly, huo2022integrating proposed an approach to integrate MPC laws of a MILP problem for a microgrid system. However, these approaches employ black-box models that do not exactly represent the solution function. Moreover, none of these works compare optimality and feasibility between their models and traditional solver benchmarks or DNN based approaches. This paper proposes a custom NN architecture that closely aligns with the PWL structure of the global solution function to improve accuracy and generalizability of mp-QP predictions.
@@ 3.1 Neural Networks
Neural Networks are universal function approximators, defined as composite functions that can be expressed as an indirect mapping between inputs x and outputs y, $f:\textbf{x}\rightarrow\textbf{y}$. The basic type of NN, known as shallow feed forward neural network, consists of an input and output layer, connected by a hidden layer, $h$. Loosely influenced by neurons and synapses of the brain, the value of each neuron is determined by the weighted sum of the previous neurons’ values, which is called input current denoted by $z_{i}$ for node $i$. However, the neuron’s value is not always defined as its input current but usually transformed by an activation function. For example, the rectified linear unit (ReLU) activation function is a threshold function that is activated only if the input current value is positive, expressed mathematically as follows:
@@
For a shallow model where $n_{h}=1$, $\textbf{W}^{1}\in\mathbb{R}^{d_{x}\times d_{h}}$, $\textbf{b}^{1}\in\mathbb{R}^{d_{h}}$$\textbf{W}^{2}\in\mathbb{R}^{d_{h}\times d_{y}}$ and $\textbf{b}^{2}\in\mathbb{R}^{d_{y}}$.
@@ 4.2 Adding Inequality Constraints
As discussed earlier, when inequality constraints are binding, the active constraints can be considered as equality constraints with the optimal dual variables, $\boldsymbol{\mu}^{*}>0$. Then, the solution can be found by expanding the Lagrangian function and finding its critical points.
@@ 4.3 Predicting Active Constraints and Shadow Prices
[…]
Proof: The reader can find the proof of the theorem in pistikopoulos2020multi .
@@
Notice that the last row of the above inverse Jacobian matrix, $\textbf{J}^{-1}_{\mathcal{B}_{i}}$, is the gradient vector $\nabla\boldsymbol{\mu}^{*}_{\mathcal{B}_{i}}$ that contains the coefficients of the linear segment of the target PWL function when a certain set of constraints, $\mathcal{B}_{i}$, are binding. In other words, the vector is the set of coefficients that maps $\textbf{C},\textbf{b}_{e},\textbf{b}_{\mathcal{B}_{i}}]$ and $\boldsymbol{\theta}$ to $\boldsymbol{\mu}^{*}_{\mathcal{B}_{i}}$.
@@ 4.4 Identification of Critical Regions and Data Generation
As explained in the Introduction, the problem characteristics in many mp-QP applications imply that most constraint combinations can never occur, so some method of filtering out the unused combinations is needed. For example, in the above 6-bus DC-OPF case with 6 inequality constraints, the set of all constraint combinations is $\{\{0\},\{1\},\{2\},\{0,1\},\dots,\{0,1,2,3,4,5\}\}$ with $2^{6}=64$ elements, yet 59 of these combinations never occur. In the power demand applications investigated in the present study, we use a commercial solver to identify the active constraint combinations by solving the problems over a range of input parameters.
begins with an input parameter, $\theta^{0}$, and solves the problem using a commercial solver at $\theta^{0}$ to identify the active set of constraints at the initial solution, $\mathcal{B}_{0}$.
Algorithm 1 details this discovery step. The algorithm first generates data starting from an initial parameter, $\boldsymbol{\theta}^{0}$ and choosing an arbitrary dimension, $k$, and incrementally increasing one dimension from zero to the highest possible feasible value. Then, after solving for the first input demand, the set of potential binding constraints, $\mathcal{B}_{0}$, is identified by checking whether the associated $\mu_{i}$ is greater than zero for the inputs. Using this active set, it expands the coefficient matrix and obtains slope coefficients $\nabla\boldsymbol{\mu}_{\mathcal{B}_{i}}$ and the matrix $J^{-1}_{\mathcal{B}_{0}}$ . This matrix is then used to calculate solutions for the input parameters and to check whether the obtained solutions satisfy the KKT conditions. If the conditions are not satisfied, a new critical region is identified, and the problem is solved again by the solver to determine the new active set, $\mathcal{B}_{i+1}$. The remaining input parameters are then solved using the updated solution function, $J^{-1}_{\mathcal{B}_{i+1}}$. This procedure is repeated until all critical regions have been discovered for all input parameters. Whenever a critical region is identified, the algorithm also provides insight into the boundaries of critical regions, $\textbf{CR}_{i}^{-},\textbf{CR}_{i}^{+}]$ corresponding to $\mathcal{B}_{i}$ that are active at solution. This information is used in Algorithm 2 to populate data without using solvers, by sampling random inputs from $\textbf{CR}_{i}^{-},\textbf{CR}_{i}^{+}]$ for all $i$ and calculating the corresponding optimal solutions using the slope corresponding to each critical region.
@@
Input: Initial point $\boldsymbol{\theta}^{0}$, increment size $\alpha$, upper limit $\theta^{+}$, KKT violation threshold toltol
@@
Solve problem for the initial point using a solver
Find active constraints: $\mathcal{B}_{0}=\{\ j\ |\ \mu_{j}>0\ \}$
Expand the coefficient matrix, calculate $\textbf{J}^{-1}_{\mathcal{B}_{0}}$ and $\nabla\boldsymbol{\mu}_{\mathcal{B}_{i}}^{*}$
Initiate $i=0$, $\textbf{CR}^{-}=[\ ]$, $\textbf{CR}^{+}=[\ ]$, $\textbf{W}^{0}=[\ ]$,
Set $\textbf{CR}^{-}[i]=\boldsymbol{\theta}$, $\textbf{CR}^{+}[i]=\boldsymbol{\theta}$ and $\textbf{W}^{0}[i]=\nabla\boldsymbol{\mu}_{\mathcal{B}_{i}}$,
@@
  Calculate $\textbf{x}^{*},\boldsymbol{\lambda}^{*},\boldsymbol{\mu}^{*}_{\mathcal{B}_{i}}]^{T}$ for $\boldsymbol{\theta}$ by substituting $\textbf{J}^{-1}_{\mathcal{B}_{i}}$ in (13)
  Check if $\textbf{x}^{*},\boldsymbol{\lambda}^{*},\boldsymbol{\mu}^{*}_{\mathcal{B}_{i}}]^{T}$ satisfies all KKT conditions in Table 1 with less than toltol error
  if KKT is not satisfied, a new critical region is found, then
   Set $i=i+1$
   (i) Solve the problem for $\boldsymbol{\theta}$ using a solver
@@
   (iii) Expand coefficient matrix to calculate $\nabla\boldsymbol{\mu}^{*}_{\mathcal{B}_{i}}$ and append to $\textbf{W}^{0}$
@@
return $\textbf{W}^{0},\textbf{CR}^{-},\textbf{CR}^{+}$